# Fair Algorithms with Probing for Multi-Agent Multi-Armed Bandits

### v5

@@ Introduction
Let $S^{*}$ be an optimal subset maximizing $\mathcal{R}(S)$. Then the set $S^{\mathrm{pr}}$ returned by Algorithm 1, for any $\zeta\geq 1$, satisfies
$\mathcal{R}(S^{\mathrm{pr}})\;\;\geq\;\frac{\,e-1\,}{\,2e-1\,}\;\frac{1}{\zeta}\mathcal{R}(S^{*}).$
@@ Acknowledgments
Let $S^{*}$ be an optimal subset maximizing $\mathcal{R}(S)$. Then the set $S^{\mathrm{pr}}$ returned by Algorithm 1, for any $\zeta\geq 1$, satisfies
$\mathcal{R}(S^{\mathrm{pr}})\;\;\geq\;\frac{\,e-1\,}{\,2e-1\,}\;\frac{1}{\zeta}\mathcal{R}(S^{*}).$
@@
Applying this to $S^{*}$, we obtain:
@@
$\leq(1-\alpha(|S^{*}|))\,f_{\mathrm{upper}}(S^{*})+h(\emptyset)$
$\leq\frac{e}{e-1}(1-\alpha(|S^{*}|))f_{\mathrm{upper}}(S_{|S^{*}|})+h(\emptyset)\quad\text{(submodular factor $1-1/e$ on $f_{\mathrm{upper}}$)}$
$\leq\frac{e}{e-1}(1-\alpha(|\widetilde{S}^{*}|))f_{\mathrm{upper}}(\widetilde{S}^{*})+h(\emptyset)$
(by selection of $\widetilde{S}^{*}$ such that $|S_{|S^{*}|}|=|S^{*}|$)
$\leq\frac{e}{e-1}(1-\alpha(|S_{j}|))f_{\mathrm{upper}}(S_{j})+h(\emptyset)$
$\quad\text{(by definition of $j=\arg\max\{(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})\}$)}.$
From the algorithm, we have that the chosen set $S^{\mathrm{pr}}$ satisfies:
$S^{\mathrm{pr}}=S_{j},$
which implies:
$\mathcal{R}(S^{\mathrm{pr}})=(1-\alpha(|S_{j}|))\mathbb{E}[\mathrm{NSW}(S_{j},\dots)].$
By the algorithm’s filtering condition,
$(1-\alpha(|S_{j}|))f_{\mathrm{upper}}(S_{j})\leq\zeta\mathcal{R}(S^{\mathrm{pr}}),$
which leads to:
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}\zeta\mathcal{R}(S^{\mathrm{pr}})+h(\emptyset).$
If $S^{\mathrm{pr}}=\emptyset$, then $\mathcal{R}(S^{\mathrm{pr}})=h(\emptyset)$ satisfies the bound trivially. Otherwise, we conclude:
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}\zeta\mathcal{R}(S^{\mathrm{pr}})+\mathcal{R}(S^{\mathrm{pr}}).$
This simplifies to:
$\mathcal{R}(S^{\mathrm{pr}})\geq\frac{e-1}{2e-1}\frac{1}{\zeta}\mathcal{R}(S^{*}).$
We assume that $\zeta$ is chosen such that Algorithm 1 always returns a non-empty probing set unless the baseline $h(\emptyset)$ is strictly better; in the latter case $S^{*}=\emptyset$ and the bound holds trivially.

### v6

@@ Introduction
Let $S^{*}$ be an optimal subset maximizing $\mathcal{R}(S)$. Let
$j^{\star}\in\arg\max_{i\in\{0,\ldots,I\}}\left\{(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})\right\}.$
Define
$\zeta_{0}\;:=\;\frac{(1-\alpha(|S_{j^{\star}}|))\,f_{\mathrm{upper}}(S_{j^{\star}})}{\mathcal{R}(S_{j^{\star}})},$
and assume $\zeta\geq\max\{1,\zeta_{0}\}$. Then the set $S^{\mathrm{pr}}$ returned by Algorithm 1 satisfies
$\mathcal{R}(S^{\mathrm{pr}})\;\geq\;\frac{e-1}{2e-1}\cdot\frac{1}{\zeta}\,\mathcal{R}(S^{*}).$
@@ Acknowledgments
Let $S^{*}$ be an optimal subset maximizing $\mathcal{R}(S)$. Let
$j^{\star}\in\arg\max_{i\in\{0,\ldots,I\}}\left\{(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})\right\}.$
Define
$\zeta_{0}\;:=\;\frac{(1-\alpha(|S_{j^{\star}}|))\,f_{\mathrm{upper}}(S_{j^{\star}})}{\mathcal{R}(S_{j^{\star}})},$
and assume $\zeta\geq\max\{1,\zeta_{0}\}$. Then the set $S^{\mathrm{pr}}$ returned by Algorithm 1 satisfies
$\mathcal{R}(S^{\mathrm{pr}})\;\geq\;\frac{e-1}{2e-1}\cdot\frac{1}{\zeta}\,\mathcal{R}(S^{*}).$
@@
Applying this inequality to the optimal set $S^{*}$ yields
@@
$\leq(1-\alpha(|S^{*}|))\,f_{\mathrm{upper}}(S^{*})+h(\emptyset).$
Since $f_{\mathrm{upper}}$ is monotone submodular, the standard greedy guarantee implies that for the greedy set $S_{|S^{*}|}$ of size $|S^{*}|$,
$f_{\mathrm{upper}}(S_{|S^{*}|})\geq\left(1-\frac{1}{e}\right)\max_{|S|=|S^{*}|}f_{\mathrm{upper}}(S)\;\;\geq\;\;\left(1-\frac{1}{e}\right)f_{\mathrm{upper}}(S^{*}),$
which is equivalent to
$f_{\mathrm{upper}}(S^{*})\leq\frac{e}{e-1}\,f_{\mathrm{upper}}(S_{|S^{*}|}).$
Substituting this into the previous bound gives
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}(1-\alpha(|S^{*}|))f_{\mathrm{upper}}(S_{|S^{*}|})+h(\emptyset).$
Let $\widetilde{S}^{*}$ be a greedy candidate with $|\widetilde{S}^{*}|=|S^{*}|$ (so that $\widetilde{S}^{*}=S_{|S^{*}|}$), hence the above inequality can be written as
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}(1-\alpha(|\widetilde{S}^{*}|))f_{\mathrm{upper}}(\widetilde{S}^{*})+h(\emptyset).$
Now define
$j^{\star}\in\arg\max_{i\in\{0,\ldots,I\}}\left\{(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})\right\}.$
By the definition of $j^{\star}$, we have
$(1-\alpha(|\widetilde{S}^{*}|))f_{\mathrm{upper}}(\widetilde{S}^{*})\leq(1-\alpha(|S_{j^{\star}}|))f_{\mathrm{upper}}(S_{j^{\star}}).$
Therefore,
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}(1-\alpha(|S_{j^{\star}}|))f_{\mathrm{upper}}(S_{j^{\star}})+h(\emptyset).$ (2)
We now relate the bound (2) to the set $S^{\mathrm{pr}}$ returned by Algorithm 1. Recall that the candidates $\{S_{i}\}_{i=0}^{I}$ are scanned in non-increasing order of $(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})$. Let $j^{\star}$ be the maximizer defined in Theorem 1. Under the assumption $\zeta\geq\zeta_{0}$, we have
$(1-\alpha(|S_{j^{\star}}|))f_{\mathrm{upper}}(S_{j^{\star}})\leq\zeta\,\mathcal{R}(S_{j^{\star}}),$
so $S_{j^{\star}}$ passes the filtering condition in Line 11. Therefore, the algorithm will return $S_{j^{\star}}$ once it is reached, unless it terminates earlier at Line 10.
Case 1: $S^{\mathrm{pr}}=\emptyset$.
If the algorithm returns $\emptyset$, it must terminate at Line 10 before accepting any set. Since $S_{j^{\star}}$ is the first candidate in the sorted order, this implies
$(1-\alpha(|S_{j^{\star}}|))f_{\mathrm{upper}}(S_{j^{\star}})<h(\emptyset).$
Because $j^{\star}$ maximizes $(1-\alpha(|S|))f_{\mathrm{upper}}(S)$, we also have
$(1-\alpha(|S_{i}|))f_{\mathrm{upper}}(S_{i})\leq(1-\alpha(|S_{j^{\star}}|))f_{\mathrm{upper}}(S_{j^{\star}})<h(\emptyset),\qquad\forall i.$
Substituting into (2) yields
$\mathcal{R}(S^{*})<\frac{e}{e-1}h(\emptyset)+h(\emptyset)=\left(\frac{e}{e-1}+1\right)h(\emptyset).$
Since $\mathcal{R}(\emptyset)=h(\emptyset)$ and $S^{\mathrm{pr}}=\emptyset$, we obtain
$\mathcal{R}(S^{\mathrm{pr}})=\mathcal{R}(\emptyset)\geq\frac{1}{\frac{e}{e-1}+1}\,\mathcal{R}(S^{*})=\frac{e-1}{2e-1}\,\mathcal{R}(S^{*})\geq\frac{e-1}{2e-1}\cdot\frac{1}{\zeta}\,\mathcal{R}(S^{*}),$
where the last inequality uses $\zeta\geq 1$.
Case 2: $S^{\mathrm{pr}}\neq\emptyset$.
In this case, by the argument above, the algorithm must return $S_{j^{\star}}$, i.e.,
$S^{\mathrm{pr}}=S_{j^{\star}}.$
Since $S^{\mathrm{pr}}$ passes Line 11, we have
$(1-\alpha(|S^{\mathrm{pr}}|))f_{\mathrm{upper}}(S^{\mathrm{pr}})\leq\zeta\,\mathcal{R}(S^{\mathrm{pr}}).$
Moreover, since the algorithm did not terminate at Line 10 when it returned $S^{\mathrm{pr}}$, we also have
$(1-\alpha(|S^{\mathrm{pr}}|))f_{\mathrm{upper}}(S^{\mathrm{pr}})\geq h(\emptyset).$
Combining the two inequalities gives
$h(\emptyset)\leq\zeta\,\mathcal{R}(S^{\mathrm{pr}}).$
Substituting $S^{\mathrm{pr}}=S_{j^{\star}}$ and the filter inequality into (2) yields
$\mathcal{R}(S^{*})\leq\frac{e}{e-1}\zeta\,\mathcal{R}(S^{\mathrm{pr}})+h(\emptyset)\leq\left(\frac{e}{e-1}+1\right)\zeta\,\mathcal{R}(S^{\mathrm{pr}}).$
Rearranging completes the proof:
$\mathcal{R}(S^{\mathrm{pr}})\geq\frac{1}{\left(\frac{e}{e-1}+1\right)\zeta}\,\mathcal{R}(S^{*})=\frac{e-1}{2e-1}\cdot\frac{1}{\zeta}\,\mathcal{R}(S^{*}).$